# Spectrum-Aware Debiasing: A Modern Inference Framework with Applications to Principal Components Regression

### v4

@@ 4.4 Asymptotic normality
We now state the asymptotic properties of the debiased PCR procedure. The proof of the theorem below is deferred to Appendix C in Appendix.
@@
Given exchangeability of entries of $\bm{\zeta}^{\star}$, we may obtain results analogous to Corollary 3.8 for finite or single coordinate inference. We defer the results to Section C.3 in Appendix.
We demonstrate Theorem 4.1 using two sets of design matrices. Our first set (top panel of Figure 3) represents more challenging variants of the settings from Figure 1. These designs contain high correlation, heterogeneity, or both. They also contain outlier eigenvalues and the signal $\bm{\beta}^{\star}$ aligns with a few top eigenvectors. Specifically, the top panel presents the following right-rotationally invariant designs: (i) $\mathsf{MatrixNormal}$-$\mathsf{B}$: stronger row- and column-wise correlations than $\mathsf{MatrixNormal}$; (ii) $\mathsf{Spiked}$-$\mathsf{B}$: larger and fewer spikes than $\mathsf{Spiked}$; (iii) $\mathsf{LNN}$-$\mathsf{B}$: matrix product with larger exponents and stronger correlations than $\mathsf{LNN}$; (iv) $\mathsf{VAR}$-$\mathsf{B}$: stronger row dependencies than $\mathsf{VAR}$; (v) $\mathsf{MultiCauchy}$: heavier tails than $\mathsf{Multi}$-$t$. Detailed description of these design distributions are given in Definition E.4 from Appendix. Our second set of experiments (bottom panel of Figure 3) uses real data designs from five domains: speech audio [42] , DNA [70], stock returns (S&P 500) [63], face images [48], and crime metrics [75]. FUrther details about these design matrices are included in Definition E.5 from Appendix.
@@ 4.5 Inference
$\mathsf{CI}_{i}\quantity(\hat{\beta}^{u}_{\mathsf{pcr},i},\hat{\tau}_{*})=\left(\hat{\beta}^{u}_{\mathsf{pcr},i}+a\sqrt{\hat{\tau}_{*}},\hat{\beta}^{u}_{\mathsf{pcr},i}+b\sqrt{\hat{\tau}_{*}}\right)$ (33)
@@ 4.6 Alignment testing
A fundamental challenge that modern data analysis presents relates to alignment of a part of the signal with eigenvectors of the sample covariance matrix. Such alignment distorts the performance of inference procedures unless they explicitly account for it. As a by-product, our Spectrum-Aware Debiasing theory provides a formal test for alignment, in other words, for testing $H_{i,0}^{\bm{\upsilon}^{\star}}:\upsilon^{\star}_{i}=0$ vs $H_{i,1}^{\bm{\upsilon}^{\star}}:\upsilon^{\star}_{i}\neq 0$, where $\upsilon^{\star}_{i}$ is given by (30). Below, $\bm{\upsilon}^{\star}$ refers to the vector with $i$-th entry $\upsilon^{\star}_{i}$. Corollary 4.3 below is proved in Section C.4 from Appendix.
@@
$\hat{\bm{\Gamma}}^{-1/2}\quantity(\hat{\bm{\theta}}_{\mathsf{pcr}}-\bm{\upsilon}^{\star})\Rightarrow\mathcal{N}(\bm{0},\mathbf{I}_{J}),$
@@
Corollary 4.3 motivates the p-values $P_{i}:=2-2\cdot\Phi\quantity(\absolutevalue{\hat{\theta}_{\mathsf{pcr},i}/s_{i}}),s_{i}:=\sqrt{\hat{\sigma}^{2}\cdot d^{-2}_{\mathcal{J}(i)}+\hat{\omega}}$. Since the quantities $\quantity(P_{i})_{i=1}^{J}$ are asymptotically independent, the Benjamini-Hochberg procedure [13] can be used to control the False Discovery Rate (FDR), which is the expected ratio of PCs falsely identified as aligned with $\bm{\beta}^{\star}$ out of all PCs identified as aligned with $\bm{\beta}^{\star}$.
@@
We demonstrate the efficacy of our alignment test in Figure 4. The top-left panel displays Benjamini-Hochberg adjusted p-values for testing alignment in the setting of the real-data designs considered in Figure 3, bottom-row. The bottom-left panel shows the true alignment angles between the underlying signal and the top six PCs. Our test accurately identifies alignment where present. In this setting, alignment detection is relatively easy, as the true alignment angles between the top PCs and signal, where present, are all significantly smaller than $90^{\circ}$ ($\angle(\bm{\beta}^{\star},\mathbf{o}_{i})\lesssim 75^{\circ}$). We illustrate in Figure 4, right-column how our alignment test performs when alignments become less pronounced and therefore harder to detect. Our method remains effective in rejecting all strong alignments present ($\angle(\bm{\beta}^{\star},\mathbf{o}_{i})\lesssim 75^{\circ}$). While it is less decisive in rejecting weak alignments ($75^{\circ}\lesssim\angle(\bm{\beta}^{\star},\mathbf{o}_{i})\lesssim 85^{\circ}$), the overall detected alignment pattern, as reflected in the small p-values shown in the top-right panel, closely matches the true alignment pattern displayed in the bottom-right panel [^3] Note that in the setting of Figure 4, we artificially aligned the signal with the 2nd, 4th and 6th PCs. However, the bottom row of Figure 4 suggests that the signal $\bm{\beta}^{\star}$ also aligns with the 1st PC for the SP500, FaceImage and Crime designs. This additional alignment was not introduced deliberately; however, it exists due to the following reason. Recall the signal decomposition from Equation 30 given by $\bm{\beta}^{\star}=\bm{\beta}^{\star}_{\mathsf{al}}+\bm{\zeta}^{\star}$. In our setting here, $\bm{\beta}^{\star}_{\mathsf{al}}$ is a linear combination of the 2nd, 4th, 6th PCs, while we generated $\bm{\zeta}^{\star}$ such that its entries have non-zero mean ($\bm{1}_{p}^{\top}\bm{\zeta}^{\star}\neq 0$). Coincidentally, the top PC of the SP500, FaceImage and Crime designs aligns with $\bm{1}_{p}$ due to correlation among the covariates. Thus, the intended alignment-complement decomposition of the signal is mis-specified, and the model is able to correctly identify the alignment of $\bm{\zeta}^{\star}$ with $\bm{1}_{p}$. We conducted similar experiments for the simulated designs in the top row of Figure 3 in Section E.4 from Appendix.
Figure 4: left-column: Same setting as Figure 3, bottom-left experiment. right-column: Same setting as Figure 3, bottom-left experiment, except that we increase difficulty of alignment detection by considering $v_{i}^{*}=\sqrt{p},i\in\{2,4,6\}$ as opposed to $v_{i}^{*}=5\cdot\sqrt{p},i\in\{2,4,6\}$. top-row: Benjamini-Hochberg adjusted p-values for $H_{i,0}^{\bm{\upsilon}^{\star}},i=1,...,6$ based on Corollary 4.3. ** indicates rejection under FDR level 0.05 and * rejection under FDR level 0.1. bottom-row: True alignment angles between the signal $\bm{\beta}^{\star}$ and $i$-th PC $\mathbf{o}_{i}$ calculated using $\mathrm{argcos}\{\mathbf{o}_{i}^{\top}\bm{\beta}^{\star}/(\norm{\mathbf{o}_{i}}_{2}.\norm{\bm{\beta}^{\star}}_{2})\}$.
5 Proof outline and novelties
Our main result relies on three main steps: (i) a characterization of the empirical distribution of a population version of $\hat{\bm{\beta}}$, (ii) connecting this population version with our data-driven Spectrum-Aware estimator, (iii) developing a consistent estimator of the asymptotic variance. We next describe our main technical novelties for step (i) in Section 5.1, and that for steps (ii) and (iii) in Section 5.2.
5.1 Result A: Distributional characterizations
Here, $\mathbf{r}_{*}$ can be interpreted as the population version of the debiased estimator $\hat{\bm{\beta}}^{u}$ and $\mathbf{r}_{**}$ as an auxiliary quantity that arises in the intermediate steps in our proof. The following theorem characterizes the empirical distribution of the entries of $\hat{\bm{\beta}}$ and $\mathbf{r}_{*}$. We prove it in Section B.1 from Appendix.
Theorem 5.1 (Distributional characterizations).
Under Assumptions 26, almost surely as $n,p\to\infty$,
$\frac{1}{p}\norm{\mathbf{y}-\mathbf{X}\hat{\bm{\beta}}}^{2}\to\tau_{**}\cdot\mathbb{E}\frac{\mathsf{D}^{2}(\eta_{*}-\gamma_{*})^{2}}{(\mathsf{D}^{2}+\eta_{*}-\gamma_{*})^{2}}+\sigma^{2}\cdot\quantity(\frac{n-p}{p}+\mathbb{E}\quantity(\frac{\eta_{*}-\gamma_{*}}{\mathsf{D}^{2}+\eta_{*}-\gamma_{*}})^{2})$
5.2 Result B: Consistent estimation of fixed points
Note that the population debiased estimator $\mathbf{r}_{*}$ cannot be used to conduct inference since $\gamma_{*}$ is unknown. Furthermore, the previous theorem says roughly that $\mathbf{r}_{*}-\bm{\beta}^{\star}$ behaves as a standard Gaussian with variance $\tau_{*}$, without providing any estimator for $\tau_{*}$. We address these two points here. In particular, we will see that addressing these points ties us to establishing consistent estimators for the solution to the fixed points defined in (20). The theorem below shows that $(\widehat{\mathsf{adj}},\hat{\eta}_{*},\hat{\tau}_{*},\hat{\tau}_{**})$ from (26) serve as consistent estimators of the fixed points $(\gamma_{*},\eta_{*},\tau_{*},\tau_{**})$, and $\hat{\bm{\beta}}^{u},\hat{\mathbf{{r}}}_{**}$ as consistent estimators of $\mathbf{r}_{*}$ and $\mathbf{r}_{**}$, where $\hat{\mathbf{{r}}}_{**}$ is defined as in (37) below. For the purpose of the discussion below, we note that $\hat{\tau}_{**}$ from (26) can be written as follows.
Note this is well-defined when
Theorem 5.2 (Consistent estimation of fixed points).
$\widehat{\mathsf{adj}}\left(p\right)\rightarrow\gamma_{*},\quad\hat{\eta}_{*}\left(p\right)\rightarrow\eta_{*},\quad\hat{\tau}_{*}\left(p\right)\rightarrow\tau_{*},\quad\hat{\tau}_{**}\left(p\right)\rightarrow\tau_{**},\quad\hat{\sigma}^{2}\left(p\right)\rightarrow\sigma^{2}$
We prove this theorem in Section B.2 from Appendix. It is not hard to see that Theorem 5.1 combined with Theorem 5.2 proves our main result Theorem 3.1.
5.3 Proof of result A
In this section, we discuss our proof novelties for Theorem 5.1. Section B.1 from Appendix contains this proof.
We use the above algorithmic proof strategy, but in case of our right-rotationally invariant designs that, the original AMP algorithms fail to apply. To alleviate this, [74] proposed vector approximate message passing algorithms. We use these algorithms to create our $\hat{\mathbf{v}}^{t}$’s. Subsequently, proving (40) presents the main challenge. To this end, one is required to show the following Cauchy convergence property of the VAMP iterates: almost surely, $\lim_{(s,t)\rightarrow\infty}\left(\lim_{p\rightarrow\infty}\frac{1}{p}\left\|\hat{\mathbf{v}}^{t}-\hat{\mathbf{v}}^{s}\right\|^{2}\right)=0.$ We prove this using a Banach contraction argument (cf. (81) from Appendix). Such an argument saw prior usage in the context of Bayes optimal learning in [59]. However, they studied a “matched” problem where the signal prior (analogous to $\mathsf{B}^{\star}$ in our setting) is known to the statistician and she uses this exact prior during the estimation process. Arguments under such matched Bayes optimal problems do not translate to our case, and proving (81) presents novel difficulties in our setting. To mitigate this, we leverage a fundamental property of the R-transform, specifically that $-zR^{\prime}(z)/R(z)<1$ for all $z$, and discover and utilize a crucial interplay of this property with the non-expansiveness of the proximal map (see Proposition A.7 (b) from Appendix).
Remark 5.1 (Comparison with [40, 41]).
In their seminal works, [40, 41] initiated the first study of the risk of regularized estimators under right-rotationally invariant designs. They stated a version of Theorem 5.1 with a partially non-rigorous argument. In their approach, an auxiliary $\ell_{2}$ penalty of sufficient magnitude is introduced to ensure contraction of AMP iterates. Later, they remove this penalty through an analytical continuation argument. However, this proof suffers two limitations. The first one relates to the non-rigorous applications of the AMP state evolution results. For instance, [41, Lemma 3] shows that for each fixed value of $p$, $\lim_{t\rightarrow\infty}\frac{\left\|\hat{x}^{t}-\hat{x}\right\|^{2}}{p}=0.$ However, in [41, Proof of Lemma 4], the authors claim that this would imply (40) upon exchanging limits with respect to $t$ and $p$. Such an exchange of limits is non-rigorous since the correctness of AMP state evolution is established for a finite number of iterations ($t<T,T$ fixed) as $p\to\infty$. The limit in $T$ is taken after $p$. The other limitation lies in the analytic continuation approach that requires multiple exchanges of limit operations [41, Appendix H] that seem difficult to justify and incur intractable assumptions [41, Assumption 1 (c), (e)] (in particular, it is unclear how to verify the existence claim in Assumption 1 (c) beyond Gaussian designs). Our alternative approach establishes contraction without the need for a sufficiently large $\ell_{2}$-regularization component, as in [40, 41], and thereby avoids the challenges associated with the analytic continuation argument.
5.4 Proof of result B
In this section, we discuss the proof of Theorem 5.2. See Section B.2 from Appendix for the proof details.
First, let us present some heuristics for how one might derive the consistent estimators $\quantity(\widehat{\mathsf{adj}},\hat{\eta}_{*},\hat{\tau}_{*},\hat{\tau}_{**})$. We start from (20a). Using Lemma 2.7, it can be written as
Recall that we have established Theorem 5.1 that shows $p\to\infty$, almost surely, $\hat{\bm{\beta}}\stackrel{{\scriptstyle W_{2}}}{{\to}}\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right).$ Combining this and (41), we expect that
Setting $\approx$ above to equality, we obtain our exact equation for the Spectrum-Aware adjustment factor, i.e.  (23). One thus expects intuitively that $\widehat{\mathsf{adj}}$ consistently estimates $\gamma_{*}$. To establish the consistency rigorously, we recognize and establish the monotonicity of the LHS of (43) as a function of $\gamma_{*}$, and study its point-wise limit. We direct the reader to Lemma B.11 and Proposition B.13 from Appendix for more details.
$\matrixquantity(\frac{1}{p}\left\|\mathbf{X}\mathbf{r}_{**}-\mathbf{y}\right\|^{2}\\ \frac{1}{p}\norm{\mathbf{y}-\mathbf{X}\hat{\mathbf{\beta}}}^{2})\approx\matrixquantity(\mathbb{E}\mathsf{D}^{2}&\delta\\ \mathbb{E}\frac{\mathsf{D}^{2}(\eta_{*}-\gamma_{*})^{2}}{(\mathsf{D}^{2}+\eta_{*}-\gamma_{*})^{2}}&\quad\frac{n-p}{p}+\mathbb{E}\quantity(\frac{\eta_{*}-\gamma_{*}}{\mathsf{D}^{2}+\eta_{*}-\gamma_{*}})^{2})\matrixquantity(\tau_{**}\\ \sigma^{2})$ (44)
@@ 6 Discussion
We conclude our paper with a discussion of two main points. First, we clarify that although our method can handle various dependencies in the design through the right-rotational invariance assumption, it does not extend to anisotropic Gaussian designs where the rows of $\mathbf{X}$ are sampled from $\mathcal{N}(\mathbf{0},\mathbf{\Sigma})$ with an arbitrary covariance matrix $\mathbf{\Sigma}$ (unless $\mathbf{\Sigma}$ is right-rotationally invariant). Moreover, contrasting with [12], our Spectrum-Aware adjustment (4) does not apply directly to non-separable penalties, e.g. SLOPE, group Lasso, etc. We note that the current framework can be expanded to address both these issues. In Appendix D from Appendix, we suggest a debiased estimator for “ellipsoidal designs” $\mathbf{X}=\mathbf{Q}^{\top}\mathbf{D}\mathbf{O}\mathbf{\Sigma}^{1/2}$ and non-separable convex penalties. We also conjecture its asymptotic normality using the non-separable VAMP formalism [37]. We leave a detailed study of this extensive class of estimators to future works.
We discuss another potential direction of extension, that of relaxing the exchangeability assumption in Corollary 3.8 and Corollary C.1 from Appendix that establish inference guarantees on finite-dimensional marginals. One may raise a related question, that of constructing confidence intervals for $\mathbf{a}^{\top}\bm{\beta}^{\star}$ for a given choice of $\mathbf{a}$. Under Gaussian design assumptions, such guarantees were obtained using the leave-one-out method as in [18, Section 4.6] or Stein’s method as in [12] without requiring the exchangeability assumption (at the cost of other assumptions on $\bm{\beta}^{\star}$ and/or $\bm{\Sigma}$). Unfortunately, these arguments no longer apply under right-rotational invariant designs owing to the presence of a global dependence structure. Thus, establishing such guarantees without exchangeability can serve as an exciting direction for future research.

### v6

@@ 4.4 Asymptotic normality
We now state the asymptotic properties of the debiased PCR procedure. The proof of the theorem below is deferred to Appendix D in Appendix.
@@
Remark 4.3.
Given exchangeability of entries of $\bm{\zeta}^{\star}$, we may obtain results analogous to Corollary 3.10 for finite or single coordinate inference. We defer the results to Section D.3 in Appendix.
Remark 4.4.
We prove that a variant of Theorem 4.1 continues to hold under a broader spectral universality class defined in Definition F.1. This result is stated in Theorem F.2.
We demonstrate Theorem 4.1 using two sets of design matrices. Our first set (top panel of Figure 3) represents more challenging variants of the settings from Figure 1. These designs contain high correlation, heterogeneity, or both. They also contain outlier eigenvalues and the signal $\bm{\beta}^{\star}$ aligns with a few top eigenvectors. Specifically, the top panel presents the following right-rotationally invariant designs: (i) $\mathsf{MatrixNormal}$-$\mathsf{B}$: stronger row- and column-wise correlations than $\mathsf{MatrixNormal}$; (ii) $\mathsf{Spiked}$-$\mathsf{B}$: larger and fewer spikes than $\mathsf{Spiked}$; (iii) $\mathsf{LNN}$-$\mathsf{B}$: matrix product with larger exponents and stronger correlations than $\mathsf{LNN}$; (iv) $\mathsf{VAR}$-$\mathsf{B}$: stronger row dependencies than $\mathsf{VAR}$; (v) $\mathsf{MultiCauchy}$: heavier tails than $\mathsf{Multi}$-$t$. Detailed description of these design distributions are given in Definition J.4 from Appendix. Our second set of experiments (bottom panel of Figure 3) uses real data designs from five domains: speech audio [50] , DNA [85], stock returns (S&P 500) [75], face images [57], and crime metrics [90]. Further details about these design matrices are included in Definition J.5 from Appendix.
@@ 4.5 Inference
[…]
@@ 4.6 Alignment testing
A fundamental challenge that modern data analysis presents relates to alignment of a part of the signal with eigenvectors of the sample covariance matrix. Such alignment distorts the performance of inference procedures unless they explicitly account for it. As a by-product, our Spectrum-Aware Debiasing theory provides a formal test for alignment, in other words, for testing $H_{i,0}^{\bm{\upsilon}^{\star}}:\upsilon^{\star}_{i}=0$ vs $H_{i,1}^{\bm{\upsilon}^{\star}}:\upsilon^{\star}_{i}\neq 0$, where $\upsilon^{\star}_{i}$ is given by (129). Below, $\bm{\upsilon}^{\star}$ refers to the vector with $i$-th entry $\upsilon^{\star}_{i}$. Corollary 4.5 below is proved in Section D.4 from Appendix.
@@
$\hat{\bm{\Gamma}}^{-1/2}\quantity(\hat{\bm{\theta}}_{\mathsf{pcr}}-\bm{\upsilon}^{\star})\Rightarrow N(\bm{0},\mathbf{I}_{J}),$
@@
[…]
@@
We demonstrate the efficacy of our alignment test in Figure 4. Panel (i) displays Benjamini-Hochberg adjusted p-values for testing alignment in the setting of the real-data designs considered in Figure 3, bottom-row. Panel (iii) shows the true alignment angles between the underlying signal and the top six PCs. Our test accurately identifies alignment where present. In this setting, alignment detection is relatively easy, as the true alignment angles between the top PCs and signal, where present, are all significantly smaller than $90^{\circ}$ ($\angle(\bm{\beta}^{\star},\mathbf{o}_{i})\lesssim 75^{\circ}$). We illustrate in Figure 4, panels (ii) and (iv) how our alignment test performs when alignments become less pronounced and therefore harder to detect. Our method remains effective in rejecting all strong alignments present ($\angle(\bm{\beta}^{\star},\mathbf{o}_{i})\lesssim 75^{\circ}$). While it is less decisive in rejecting weak alignments ($75^{\circ}\lesssim\angle(\bm{\beta}^{\star},\mathbf{o}_{i})\lesssim 85^{\circ}$), the overall detected alignment pattern, as reflected in the small p-values shown in panel (ii), closely matches the true alignment pattern displayed in panel (iv) [^5] Note that in the setting of Figure 4, we artificially aligned the signal with the 2nd, 4th and 6th PCs. However, the bottom row of Figure 4 suggests that the signal $\bm{\beta}^{\star}$ also aligns with the 1st PC for the SP500, FaceImage and Crime designs. This additional alignment was not introduced deliberately; however, it exists due to the following reason. Recall the signal decomposition from Equation 129 given by $\bm{\beta}^{\star}=\bm{\beta}^{\star}_{\mathsf{al}}+\bm{\zeta}^{\star}$. In our setting here, $\bm{\beta}^{\star}_{\mathsf{al}}$ is a linear combination of the 2nd, 4th, 6th PCs, while we generated $\bm{\zeta}^{\star}$ such that its entries have non-zero mean ($\bm{1}_{p}^{\top}\bm{\zeta}^{\star}\neq 0$). Coincidentally, the top PC of the SP500, FaceImage and Crime designs aligns with $\bm{1}_{p}$ due to correlation among the covariates. Thus, the intended alignment-complement decomposition of the signal is mis-specified, and the model is able to correctly identify the alignment of $\bm{\zeta}^{\star}$ with $\bm{1}_{p}$.. We conducted similar experiments for the simulated designs in the top row of Figure 3 in Section J.4 from Appendix.
Figure 4: Panels (i) and (iii) use the Figure 3 (bottom-left) setting with $v_{i}^{\star}=5\sqrt{p}$ for $i\in\{2,4,6\}$; (ii) and (iv) weaken the alignment to $v_{i}^{\star}=\sqrt{p}$. Panels (i)(ii) report Benjamini–Hochberg–adjusted $p$-values for $H^{\alpha^{\star}}_{i,0}$ via Corollary 4.5 (**: FDR 0.05, *: 0.1). Panels (iii)(iv) show the corresponding true alignment angles $\angle(\mathbf{o}_{i},\bm{\beta}^{\star})$.
@@ 5 Discussion
We conclude our paper with a discussion of two main points. First, we clarify that although our method can handle various dependencies in the design through the right-rotational invariance assumption, it does not extend to anisotropic Gaussian designs where the rows of $\mathbf{X}$ are sampled from $N(\mathbf{0},\mathbf{\Sigma})$ with an arbitrary covariance matrix $\mathbf{\Sigma}$ (unless $\mathbf{\Sigma}$ is right-rotationally invariant). Moreover, contrasting with [13], our Spectrum-Aware adjustment (4) does not apply directly to non-separable penalties, e.g. SLOPE, group Lasso, etc. We note that the current framework can be expanded to address both these issues. In Appendix I from Appendix, we suggest a debiased estimator for “ellipsoidal designs” $\mathbf{X}=\mathbf{Q}^{\top}\mathbf{D}\mathbf{O}\mathbf{\Sigma}^{1/2}$ and non-separable convex penalties. We also conjecture its asymptotic normality using the non-separable VAMP formalism [44]. We leave a detailed study of this extensive class of estimators to future works.
We discuss another potential direction of extension, that of relaxing the exchangeability assumption in Corollary 3.10 and Corollary D.1 from Appendix that establish inference guarantees on finite-dimensional marginals. One may raise a related question, that of constructing confidence intervals for $\mathbf{a}^{\top}\bm{\beta}^{\star}$ for a given choice of $\mathbf{a}$. Under Gaussian design assumptions, such guarantees were obtained using the leave-one-out method as in [20, Section 4.6] or Stein’s method as in [13] without requiring the exchangeability assumption (at the cost of other assumptions on $\bm{\beta}^{\star}$ and/or $\bm{\Sigma}$). Unfortunately, these arguments no longer apply under right-rotational invariant designs owing to the presence of a global dependence structure. Thus, establishing such guarantees without exchangeability can serve as an exciting direction for future research.
@@
SUPPLEMENTARY MATERIAL: “SPECTRUM–AWARE DEBIASING: A MODERN INFERENCE FRAMEWORK
WITH APPLICATIONS TO PRINCIPAL COMPONENTS REGRESSION”
By Yufan Li1,a  and  Pragya Sur1,b
1Department of Statistics, Harvard University   ayufan_li@g.harvard.edu; bpragya@fas.harvard.edu